OpenStudy (anonymous):

If it takes 11/2 hours to fill an aquarium and 11/3 hours to drain it, how long will it take to empty the tank and refill it

OpenStudy (e.mccormick):

OK. What did you have trouble with on this?

OpenStudy (anonymous):

i added 11/3 and 11/2 which gave me 55/6

OpenStudy (anonymous):

how do i convert it to hour

OpenStudy (e.mccormick):

It is still hours. No conversion. 55/6 hours.

OpenStudy (e.mccormick):

They gave you two number in hours so you got an answer in hours. Now, if they gave you one in hours and one in minutes, you would need to convert first to they were both in the same units before adding. But, because they gave it to you all in hours, no problems.

OpenStudy (anonymous):

the options are a)2 hours and 3 min b) 2 hours and 40 mins c) 2 hours and 45 mins d) 2 hours and 50 mins e) 2 hours and 55 mins

OpenStudy (anonymous):

i mean a) 2 hours and 30 mins

OpenStudy (e.mccormick):

AAAH! Ok.Is that 11/2 and 11/3 or is it 1 and 1/2, and 1 and 1/3?

OpenStudy (anonymous):

it's 11/2 and 11/3

OpenStudy (mertsj):

Perhaps it sis \[1\frac{1}{2}\]

OpenStudy (e.mccormick):

Yah, which is why I was asking because that would make it add to one of the answers.

OpenStudy (mertsj):

Exactly. If it is \[\frac{11}{2}\] the answer is not there.

OpenStudy (e.mccormick):

\[1\frac{1}{2}+1\frac{1}{3}=?\]

OpenStudy (e.mccormick):

If you look at this as 1 hour plus half an hour plus one hour plus a third of an hour, it becomes two hours and? Half and hour and a third of an hour are how many minutes?

OpenStudy (anonymous):

i have one more question

OpenStudy (anonymous):

If 5/7 of the lawn is cut in 35 mins, then how much more time, in minutes, will it take to finish the job

OpenStudy (e.mccormick):

OK. So This is a ratio problem with a twist. Do you know how to set those up?

OpenStudy (anonymous):

no

OpenStudy (e.mccormick):

OK. The big thing here is you have 5/7 of a lawn but you want a whole lawn. A whole lawn is one lawn. So, 5/7 + x = 1. What is x? Lets start there.

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

So, can you solve that? What added to 5/7 makes for 1, or another way to say that is:\[\frac{5}{7}+\frac{?}{?}=\frac{7}{7}\]

OpenStudy (anonymous):

7/7- 5/7

OpenStudy (e.mccormick):

Yes. Which is?

OpenStudy (anonymous):

2/7

OpenStudy (e.mccormick):

OK. Now you know how much of the lawn is the rest of the lawn. So lets use L for lawn.\[\frac{5}{7}L=35\]and what you want is \[\frac{2}{5}L=Answer\] There are a couple ways to solve it at that point. Can you think of one that might work?

OpenStudy (e.mccormick):

Sorry. Second is supose to be 2/7L Oops!

OpenStudy (anonymous):

no

OpenStudy (e.mccormick):

OK. Lets start by finding what those 5/7ths are. So you know you have 35, and that represents 5 of something broken into 7 parts. To me, that means 5 of something is somehow related to 35. So what do you get if you divide 35 by 5?

OpenStudy (e.mccormick):

Mathematically this is: \[\frac{5}{7}L*\frac{1}{5}=\frac{35}{1}*\frac{1}{5}\]

OpenStudy (e.mccormick):

Now, if you simplify that, what would you get on the right and left. You can ignore the L for doing the simplification and just focus on the fractions.

OpenStudy (e.mccormick):

Well, it says you are offline. Once you know that simolified form above, you will have 1/7th L = something. You want 2/7ths, so multiply the something by 2 and you have the answer.

OpenStudy (anonymous):

I got it. thank you very much.